# Tuning-free Quasi-stiffness Control Framework of a Powered Transfemoral Prosthesis for Task-adaptive Walking

### v2

@@ Abstract
Impedance-based control represents a prevalent strategy in the development of powered transfemoral prostheses. However, creating a task-adaptive, tuning-free controller that effectively generalizes across diverse locomotion modes and terrain conditions continues to be a significant challenge. This letter proposes a tuning-free and task-adaptive quasi-stiffness control framework for powered prostheses that generalizes across various walking tasks, including the torque-angle relationship reconstruction part and the quasi-stiffness controller design part. A Gaussian Process Regression (GPR) model is introduced to predict the target features of the human joint’s angle and torque in a new task. Subsequently, a Kernelized Movement Primitives (KMP) is employed to reconstruct the torque-angle relationship of the new task from multiple human reference trajectories and estimated target features. Based on the torque-angle relationship of the new task, a quasi-stiffness control approach is designed for a powered prosthesis. Finally, the proposed framework is validated through practical examples, including varying speeds and inclines walking tasks. Notably, the proposed framework not only aligns with but frequently surpasses the performance of a benchmark finite state machine impedance controller (FSMIC) without necessitating manual impedance tuning and has the potential to expand to variable walking tasks in daily life for the transfemoral amputees.
@@
Individuals with transfemoral amputations suffer difficulties in ambulation, particularly during demanding locomotion activities such as ascending ramps and stairs. Powered transfemoral prostheses hold promise in restoring the lost joint function and significantly enhancing the mobility of these individuals [1, 2]. While significant progress has been made in both the mechanical designs [3, 4, 5, 6, 7, 8] and controller designs [9, 10, 11, 12, 13, 14, 15, 16] of powered prostheses over the past y, developing a prosthesis control framework that generalizes across various locomotion modes and terrain conditions remains a challenge [2].
@@
The main contributions are summarized as follows:
@@ II-A Overview of the Framework
As shown in Fig. 1, the proposed tuning-free, task-adaptive quasi-stiffness control framework consists of a task estimation part, a torque-angle relationship reconstruction part, and a quasi-stiffness control part. The task estimation part perceives amputee locomotion information and terrain conditions. It provides task parameters ( where denotes walking speed and represents terrain parameters, such as inclination and stair height .) for the next part. Previous research, including our own, has explored a variety of task estimation methodologies, encompassing vision-based approaches [33, 34] for assessing terrain conditions and IMU (Inertial Measurement Unit)-based strategies [19, 35, 36] for estimating walking speed. These methodologies facilitate the estimation of task parameters .
@@ II-B Target Features Prediction
An open-source dataset [37] is implemented as the human reference trajectories, which contains tasks including multiple locomotion modes (flat walking, ramp climbing, and stair climbing) and terrain conditions (walking speeds, ramp inclinations, and stair heights). To obtain the new task-specific torque-angle relationship, this study introduces GPR-based learning from the reference trajectory method for target features prediction, as shown in Fig. 1(b). Assume there are reference trajectories and define the train data as
@@ III-A Experimental Setup and Protocol
2) Experimental Protocol: A transfemoral amputee patient (44 years old, 60 kg, 170 cm, 24 years post-amputation) is invited to participate in the experiments to evaluate the proposed learning quasi-stiffness control framework. To investigate the task-adaptive performance of the proposed control framework, steady-state walking task trails including varying speeds (from 0.4 m/s to 1.0 m/s, m/s deviations) and varying inclines (from 2 deg to 8 deg, deg deviations) are conducted on an instrumented split-belt treadmill (Bertec, Columbus, OH, USA). The initial trials began with a period of acclimation in which the amputee participant first walked on the treadmill at a self-preferred speed (about 0.6 m/s) until feeling comfortable. After tuning and acclimation for a period of time, the participant walked on the treadmill again and performed each steady-state task trail for 30 seconds. The participant provided informed consent, and the Southern University of Science and Technology Institutional Review Board approved the experiments.
@@ III-B Experimental Results
1) Target Features Estimation: Fig. 1(b) shows the experimental results of target features prediction, where the blue dots denote the reference target features and the purple dots represent the predicted target features of a new task (new speed). The GPR-based target features prediction model is imported into the embedded platform of the powered prosthesis and used for feature prediction of new tasks.
2) Torque-angle Relationship Reconstruction: Fig. 4 shows the torque-angle relationship learning and reconstruction in a new task. The distribution of the human reference torque-angle relationship can be learned by GMM. Based on the learned GMM, the reference torque-angle relationship can be reproduced by GMR. As shown in Fig. 4(a), the blue torque-angle curves denote the human reference trajectoires of different speed walking tasks, and the yellow torque-angle curve represents the reference torque-angle relationship reproduced by GMR. As shown in Fig. 4(b), the yellow dashed curve represents a torque-angle relationship reconstructed by the human reference torque-angle relationships (blue curves) and target features (purple dots) in a new task.
@@
4) Amputee Participant Experimental Results: Fig. 6 shows a series of gait trails of a transfemoral amputee performing varying speed and inclines walking tasks. Fig. 7 presents that the kinematic and kinetics trajectories produced by the proposed method are similar to the able-bodied references. The powered prosthetic joint trajectories show strong similarities to the intact joint trajectories, which also indicate that the quai-stiffness can improve gait symmetry. Fig. 8 shows the root mean squared error (RMSE) in the powered prosthetic joint trajectories relative to able-bodied trajectories and intact leg trajectories during the walking experiments. The lower RMSE of the observed kinematics with respect to the intact leg kinematics indicates that the proposed quasi-stiffness control achieves better joint kinematic symmetry. For comparison, Fig. 9 shows the observed mean kinematics and kinetics trajectories of the proposed learning quasi-stiffness control framework and the benchmark FSMIC. Fig. 10 shows the RMSE of the observed kinematics and kinetics generated by the proposed method and the benchmark FSMIC with respect to the able-bodied references. The angle and moment RMSE of the proposed method are usually smaller than the benchmark.
@@ IV-A Evaluation of the Proposed Framework
This work proposed a learning task-adaptive quasi-stiffness controller for a powered transfemoral prosthesis that generalizes across tasks. The proposed method has the potential to quickly and autonomously configure the impedance parameters for variable tasks. Experimental results and walking performance (see the accompanying video) highlight its capacity to replicate able-bodied biomechanics across variable tasks. As shown in Fig. 7, the observed kinematic and kinetic profiles of the powered prosthetic joint vary with walking tasks and are similar to able-bodied references, such as appropriately varying peak angles, peak moments, and trends of profiles. As speed or incline increases, the proposed method can generate biomechanical performances similar to the intact limb of the user and healthy individuals. For instance, as the incline increases, ankle joint torque correspondingly increases to provide individuals with limb amputations more push-off torque. The powered prosthetic joint trajectories show strong similarities to the intact joint trajectories (also the lower RMSE of the observed kinematics with respect to the intact leg kinematics shown in Fig. 8(a)), which indicates that the quasi-stiffness control has the potential to improve gait symmetry. Knee kinematics show the most differences with respect to able-bodied data, which may be because the torque-angle relationship of the knee joint is more complex with fewer linear parts, which inevitably introduces errors when using linear approximations.
This article introduces the widely used FSMIC controller as a benchmark for comparison. Fig. 9 and 10 show the proposed learning quasi-stiffness control framework meets the benchmark’s performance in most walking tasks. Fig. 10(b) indicates that the proposed method surpasses the benchmark’s performance in generating the biological kinetics. As shown in Fig. 9 and Fig. 10(a), the FSMIC’s performance was not drastically worse than the proposed method in the tested metrics, it required an average of 20 min of tuning per tuned task. The proposed method highlights its advantages in reproducing biomimetic kinematics and kinetics similar to able-bodied references, aligns with the performance of the FSMIC without requiring manual impedance tuning, and reveals the potential to enable autonomous task adaptation during continuously varying walking tasks.

### v3

# A Learning Quasi-stiffness Control Framework of a Powered Transfemoral Prosthesis for Adaptive Speed and Incline Walking

@@ Abstract
Impedance-based control represents a prevalent strategy in the powered transfemoral prostheses because of its ability to reproduce natural walking. However, most existing studies have developed impedance-based prosthesis controllers for specific tasks, while creating a task-adaptive controller for variable-task walking continues to be a significant challenge. This article proposes a task-adaptive quasi-stiffness control framework for powered prostheses that generalizes across various walking tasks, including the torque-angle relationship reconstruction part and the quasi-stiffness controller design part. A Gaussian Process Regression (GPR) model is introduced to predict the target features of the human joint’s angle and torque in a new task. Subsequently, a Kernelized Movement Primitives (KMP) is employed to reconstruct the torque-angle relationship of the new task from multiple human reference trajectories and estimated target features. Based on the torque-angle relationship of the new task, a quasi-stiffness control approach is designed for a powered prosthesis. Finally, the proposed framework is validated through practical examples, including varying speeds and inclines walking tasks. Notably, the proposed framework not only aligns with but frequently surpasses the performance of a benchmark finite state machine impedance controller (FSMIC) without necessitating manual impedance tuning and has the potential to expand to variable walking tasks in daily life for the transfemoral amputees.
@@
Individuals with transfemoral amputations suffer difficulties in ambulation, particularly during demanding locomotion activities such as ascending ramps and stairs. Powered transfemoral prostheses hold promise in restoring the lost joint function and significantly enhancing the mobility of these individuals [1, 2]. While significant progress has been made in both the mechanical designs [3, 4, 5, 6, 7, 8] and controller designs [9, 10, 11, 12, 13, 14, 15, 16] of powered prostheses over the past years, developing a prosthesis control framework that generalizes across various locomotion modes and terrain conditions remains a challenge [2].
@@
This study aims to develop a task-adaptive, tuning free quasi-stiffness control framework for a powered transfemoral prosthesis that can reproduce natural walking and biomimic joint kinematics and kinetics similar to those of able-bodied people. The main contributions are summarized as follows:
@@ II-A Overview of the Framework
As shown in Fig. 1, the proposed task-adaptive quasi-stiffness control framework consists of a task estimation part, a torque-angle relationship reconstruction part, and a quasi-stiffness control part. The task estimation part perceives amputee locomotion information and terrain conditions. It provides task parameters ( where denotes walking speed and represents terrain parameters, such as inclination and stair height .) for the next part. Previous research, including our own, has explored a variety of task estimation methodologies, encompassing vision-based approaches [33, 34] for assessing terrain conditions and IMU (Inertial Measurement Unit)-based strategies [19, 35, 36] for estimating walking speed. These methodologies facilitate the estimation of task parameters . In this study, we used the speed estimation method in [19, 36] and the incline estimation method in [33, 37]. The task estimate accuracy is shown in Table I.
Task | Accuracy | Reference Method |
Speed | 94 % | [19, 36] |
Incline | 95 % | [33, 37] |
@@ II-B Target Features Prediction
Two open-source dataset [38, 39] are implemented as the human reference trajectories, which contains tasks including multiple locomotion modes (flat walking and ramp climbing) and terrain conditions (walking speeds (28 speeds ranging from 0.5 to 1.85 m/s in 0.05 m/s increments) and ramp inclinations (6 inclination angles in [38] and 4 inclination angle in [39]), rang from -10 deg to 10 deg with 5 deg increments). Notably, for varying walking speeds on treadmill, we divided 70 % of the data as training set and 30 % as testing set; for varying inclines, we used all the incline walking data in [38] and [39] as training set and the data we collected under the motion capture system (rang from -8 deg to 8 deg in 2 deg increments) as testing set. To obtain the new task-specific torque-angle relationship, this study introduces GPR-based learning from the reference trajectory method for target features prediction, as shown in Fig. 2(b). Assume there are reference trajectories and define the train data as
@@ III-A Experimental Setup and Protocol
2) Experimental Protocol: A transfemoral amputee patient (44 years old, 60 kg, 170 cm, 24 years post-amputation) is invited to participate in the experiments to evaluate the proposed learning quasi-stiffness control framework. Notably, to evaluate the performance of the GPR model used in the target features estimation and the KMP model used in the torque-angle curve reconstruction, we used the walking speeds (from 0.2m/s to 0.8 m/s in 0.2 m/s increments) and ramp inclinations (from -8 deg to 8 deg in 2 deg increments) as test set. To investigate the task-adaptive performance of the proposed control framework, steady-state walking task trails including varying speeds (from 0.4 m/s to 1.0 m/s, m/s deviations) and varying inclines (from 2 deg to 8 deg, deg deviations) are conducted on an instrumented split-belt treadmill (Bertec, Columbus, OH, USA). The initial trials began with a period of acclimation in which the amputee participant first walked on the treadmill at a self-preferred speed (about 0.6 m/s) until feeling comfortable. After tuning and acclimation for a period of time, the participant walked on the treadmill again and performed each steady-state task trail for 30 seconds. The participant provided informed consent, and the Southern University of Science and Technology Institutional Review Board approved the experiments.
@@ III-B Experimental Results
1) Target Features Estimation: Fig. 2(b) shows the experimental results of target features prediction, where the blue dots denote the reference target features and the purple dots represent the predicted target features of a new task (new speed). The GPR-based target features prediction model is imported into the embedded platform of the powered prosthesis and used for feature prediction of new tasks. The prediction errors were calculated by RMSE (root mean squared error) between reference and predicted target joint angles and joint torques, as shown in Fig. 4.
2) Torque-angle Relationship Reconstruction: Fig. 5 shows the torque-angle relationship learning and reconstruction in a new task. The distribution of the human reference torque-angle relationship can be learned by GMM. Based on the learned GMM, the reference torque-angle relationship can be reproduced by GMR. As shown in Fig. 5(a), the blue torque-angle curves denote the human reference trajectoires of different speed walking tasks, and the yellow torque-angle curve represents the reference torque-angle relationship reproduced by GMR. As shown in Fig. 5(b), the yellow dashed curve represents a torque-angle relationship reconstructed by the human reference torque-angle relationships (blue curves) and target features (purple dots) in a new task. The reconstruction errors were also calculated by RMSE between reference and predicted torque-angle relationship, as shown in Fig. 6.
@@
4) Amputee Participant Experimental Results: Fig. 8 shows a series of gait trails of a transfemoral amputee performing varying speed and inclines walking tasks. Fig. 9 and 10 present that the kinematic and kinetics trajectories produced by the proposed method are similar to the able-bodied references. The powered prosthetic joint trajectories show strong similarities to the intact joint trajectories, which also indicate that the quai-stiffness can improve gait symmetry. Fig. 11 shows the RMSE in the powered prosthetic joint trajectories relative to able-bodied trajectories and intact leg trajectories during the walking experiments. The lower RMSE of the observed kinematics with respect to the intact leg kinematics indicates that the proposed quasi-stiffness control achieves better joint kinematic symmetry. For comparison, Fig. 12 shows the observed mean kinematics and kinetics trajectories of the proposed learning quasi-stiffness control framework and the benchmark FSMIC. Fig. 13 shows the RMSE of the observed kinematics and kinetics generated by the proposed method and the benchmark FSMIC with respect to the able-bodied references. The angle and moment RMSE of the proposed method are usually smaller than the benchmark.
@@ IV-A Evaluation of the Proposed Framework
This work proposed a learning task-adaptive quasi-stiffness controller for a powered transfemoral prosthesis that generalizes across tasks. The proposed method has the potential to quickly and autonomously configure the impedance parameters for variable tasks. Experimental results and walking performance (see the accompanying video) highlight its capacity to replicate able-bodied biomechanics across variable tasks. As shown in Fig. 9 and 10, the observed kinematic and kinetic profiles of the powered prosthetic joint vary with walking tasks and are similar to able-bodied references, such as appropriately varying peak angles, peak moments, and trends of profiles. As speed or incline increases, the proposed method can generate biomechanical performances similar to the intact limb of the user and healthy individuals. For instance, as the incline increases, ankle joint torque correspondingly increases to provide individuals with limb amputations more push-off torque. The powered prosthetic joint trajectories show strong similarities to the intact joint trajectories (also the lower RMSE of the observed kinematics with respect to the intact leg kinematics shown in Fig. 11(a)), which indicates that the quasi-stiffness control has the potential to improve gait symmetry. Knee kinematics show the most differences with respect to able-bodied data, which may be because the torque-angle relationship of the knee joint is more complex with fewer linear parts, which inevitably introduces errors when using linear approximations.
This article introduces the widely used FSMIC controller as a benchmark for comparison. Fig. 12 and 13 show the proposed learning quasi-stiffness control framework meets the benchmark’s performance in most walking tasks. Fig. 13(b) indicates that the proposed method surpasses the benchmark’s performance in generating the biological kinetics. As shown in Fig. 12 and Fig. 13(a), the FSMIC’s performance was not drastically worse than the proposed method in the tested metrics, it required an average of 20 min of tuning per tuned task. The proposed method highlights its advantages in reproducing biomimetic kinematics and kinetics similar to able-bodied references, aligns with the performance of the FSMIC without requiring manual impedance tuning, and reveals the potential to enable autonomous task adaptation during continuously varying walking tasks.